Verify that y=sin 3t+2cos 3t is a solution to the initial value problem 2y +18y = 0; y(0) = 2, y (O) = 3. Find the derivat
Answer #1

y=sinat +2 CB 3t 2y +189 = 0 Y10)=2, y(0) -3 Salution: Y = sinst +220334 > 0 Differente iating 0 evorit t, we get ys at INOW, : . Y (0) = 2 y(0) - 303 (0) - 6 Sino = 3.1–6.0 yo) 3 into ay+18y = o gives ... Intersection Sin3t+2 cos 37 -0, 2 +18